together, Thabo and his sister are twenty four years old. Three years ago, Thabo was twice as old as his sister was then. how old are they now ?
Found 2 solutions by Theo, ikleyn:
Answer by Theo(13342)

let x = thabo's age now.
let y = his sister's age now.

you are given that x + y = 24
you are also given that (x - 3) = 2 * (y - 3)

simplify the second equation to get:
x - 3 = 2y - 6
add 3 to both sides of the second equation and subtract y from both sides of it to get:
x - 2y = 3

your two equations that need to be solved simultaneously are:

x + y = 24
x - 2y = 3

subtract the second equation from the first to get:
3y = 27
solve for y to get:
y = 9

note that x - x = 0
note also that y - (-2y) = y + 2y = 3y

since x + y = 24, then x must be equal to 15 because 15 + 9 = 24

confirm the solution is correct by replacing x with 15 and y with 9 in the original equations.

x + y = 24 becomes 15 + 9 = 24 which becomes 24 = 14 which is correct.
(x - 3) = 2 * (y - 3) becomes (15 - 3) = 2 * (9 - 3) which becomes 12 = 12 which is also correct.

the solution is confirmed to be good.
the solution is that that thabo is cureently 15 years old and his sister is 9.



Answer by ikleyn(53765)

.
together, Thabo and his sister are twenty four years old.
Three years ago, Thabo was twice as old as his sister was then.
how old are they now ?
~~~~~~~~~~~~

Three years ago, the sum of their ages was, OBVIOUSLY, 24 - 3 - 3 = 18 years,

AND Thabo was twice as old as his sister.


At this point, it is OBVIOUS, that three years ago Thabo was 12 years old, and the sister was 6 years old.


Hence, now Thabo is 12+3 = 15 years old, while his sister is 6+3 = 9 years old.    ANSWER

----------

So,  the problem is easily solved  MENTALLY,  using logical reasoning.